OpenStudy (vane11):

A rectangular box is to have a square base and a volume of 20 ft3. If the material for the base costs $0.35 per square foot, the material for the sides costs $0.10 per square foot, and the material for the top costs $0.15 per square foot, determine the dimensions of the box that can be constructed at minimum cost.

OpenStudy (zzr0ck3r):

let the length of the base be x let the height be h then the volume \(V= l*w*h = x^2h\) the are of the top and bottom of the base are each is x^2 the are of the 4 sides is 4xh so all together we get cost \(C=.35x^2+.15x^2+.1*4xh\) we know h = V/x^2 = 20x^2 plug that in \(C = .35x^2+.15x^2+.1*4x(20/x^2) = .35x^2+.15x^2+.1*4(20/x)\)

OpenStudy (zzr0ck3r):

can you handle it from here?

OpenStudy (zzr0ck3r):

@Vane11 ?

OpenStudy (zzr0ck3r):

that should say we know h = V/x^2 = 20/x^2

OpenStudy (vane11):

I think I can let me try and take it from here, I'll let you know if I get stuck, thanks!

OpenStudy (vane11):

got it, 2, 2, and 5, Thanks again!

OpenStudy (zzr0ck3r):

sounds right